Circular No. 7331

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1999 X1 (HUG-BELL) Gary Hug and Graham E. Bell, Eskridge, KS, report their discovery of

a comet (discovery observation given below) showing a faint tail in p.a. 285^{o} on CCD images taken with a 0.3-m Schmidt-Cassegrain reflector on Dec. 10. Following posting on the NEO Confirmation Page, L. Šarounová (Ondřejov, 0.65-m reflector) obtained observations on Dec. 11.2 UT showing a tail 20′′ long in p.a. ∼ 300^{o}. C. Hergenrother, Lunar and Planetary Laboratory, reports that a co-added 1200-s R-band image obtained with the 1.54-m Kuiper telescope on Dec. 11 shows a 15′′ coma and a slightly curved tail > 1′ long in p.a. 280^{o}.

1999 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 10.33347  8^{h}24^{m}32.^{s}66  +20^{o}07′54′′.0  18.8

All of the available astrometry (including prediscovery observations on Oct. 10 and Dec. 7 by LINEAR), along with the following orbital elements, appear on MPEC 1999-X24:

T = 1999 June 20.7057 TT  ω = 296.^{o}8671
e = 0.471559  Ω = 103.6425 }2000.0
q = 1.936291 AU  i = 10.9682
a = 3.664160 AU n^{o} = 0.1405214 P = 7.014 years

SUPERNOVA 1999ge IN NGC 309 S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999ge, obtained by A. Mahdavi on Dec. 7.13 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), indicates it to be a type-II supernova at an early epoch. The spectrum exhibits a generally featureless and very blue continuum. The most prominent feature is weak Hδ absorption, though absorption from Hβ and Hγ is also visible, as well as a hint of a P-Cyg Hα profile. Cross-correlation of the host-galaxy spectrum with an emission-line template yields a recession velocity of 5650 ± 30 km/s for NGC 309.

SUPERNOVA 1999em IN NGC 1637 CCD photometry by J. Safar and D. Hanžl, Brno, Czech Republic:

Nov. 2.051 UT, V = 13.89 ± 0.03; 15.943, V = 13.92 ± 0.01, B − V = +0.75 ± 0.2; Dec. 2.903, V = 14.00 ± 0.05.

1999 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green